two dice are rolled. Find the probability that the score on the dice is either 7 or 9
Answer by stanbon(75887)

two dice are rolled. Find the probability that the score on the dice is either 7: 1/6 ; 2*5 ; 3/4 ; 4/3 ; 5/2 ; 6/1
------
9: 3/6 ; 4/5 ; 5/4 ; 6/3
-------
P(7 or 9) = (6/36) + (4/36) = 10/36 = 5/18
============================================
Cheers,
Stan H.
=====